# Fractionalization as an alternate to charge ordering in electronic insulators

### v1

@@ C.2 Time reversal invariant Abelian TOs
and thus $0=2\theta_{a}=\theta_{a,a}$ modulo $2\pi$. We will thus have that $\theta_{a,a}=0$ for every $a\in\mathcal{C}=\mathrm{Ker}(g)/\mathrm{Im}(f)$. Now we note that $02\theta_{a,b}=\theta_{a^{2},b}=$ 0 since $a\in\mathcal{C}$ has order two. Thus $\theta_{a,b}=0,\pi$ for every $a,b\in\mathcal{C}$.
Now take some $a\in\mathcal{C}$, then $\theta_{a,a}=0$ as demonstrated above. Since braiding is non-degenerate on $\mathcal{C}$ there must be some $b\in\mathcal{C}$ such that $\theta_{a,b}=\pi$. Then take $\mathcal{S}=\{1,a,b,ab\}\subseteq\mathcal{C}$. Then by Thm. 21 we see that $|\mathcal{C}|=|\mathcal{S}||\mathcal{S}^{\perp}|=4|\mathcal{S}^{\perp}|$. We further know by Corr. C.9 that $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$, so the only elements which braid trivially with everything in $\mathcal{S}^{\perp}$ lie in $\mathcal{S}$. But every nontrivial element of $\mathcal{S}$ braids nontrivially with at least one other element of $\mathcal{S}$ and thus $\mathcal{S}\cap\mathcal{S}^{\perp}=\{1\}$. We conclude that braiding is non-degenerate when restricted to $\mathcal{S}^{\perp}$. Thus $\mathcal{S}^{\perp}\subseteq\mathcal{C}$ is itself a bosonic TO. We can then find some $\mathcal{S}^{\prime}=\{1,a^{\prime},b^{\prime},a^{\prime}b^{\prime}\}\subseteq\mathcal{S}^{\perp}$ where $\theta_{a^{\prime},b^{\prime}}=\pi$ and again apply Thm. 21 to see that $|\mathcal{S}^{\perp}|=4|\mathcal{S}^{\prime\perp}|$, where $\mathcal{S}^{\prime\perp}\subseteq\mathcal{S}^{\perp}$ is another bosonic TO. Continuing to apply this logic until we have exhausted the finite set $\mathcal{C}$ reveals that $|\mathcal{C}|=4^{m}=2^{2m}$ for some $m$. Thus $|\mathcal{C}|$ is a square and we see $|\mathcal{A}_{b}|=|\mathrm{Im}(f)|^{2}|\mathcal{C}|$ must also be a square. ∎
@@
If $\mathcal{A}$ is a time-reversal invariant fermionic Abelian TO then we write $\mathcal{A}_{b}=\mathcal{A}/\{1,c\}$. Braiding will become non-degenerate on this TO. This is thus a bosonic TO with a $\pi$-ambiguity in the topological spin, since we have identified the fundamental fermion $c$ with the vacuum. However, there is no ambiguity of braiding.
We can now repeat the arguments of Thm. C.19 to see that $|\mathcal{A}_{b}|=|\mathrm{Im}(f)|^{2}|\mathcal{C}|$, where $\mathcal{C}=\mathrm{Ker}(g)/\mathrm{Im}(f)$. We continue to have that every element of $\mathcal{C}$ is order two, so $\mathcal{C}=\mathbb{Z}_{2}^{n}$ for some $n$. All of these arguments did not rely on topological spin, only on braiding. The first result then follows immediately. We know that $|\mathcal{A}|=2|\mathcal{A}_{b}|=|\mathrm{Im}(f)|^{2}2^{n+1}$. If $n$ is odd then $|\mathcal{A}|=l^{2}$ for some $l$, while if $n$ is even then $|\mathcal{A}|=2l^{2}$ for some $l$.
We will need to work a little harder to prove the second result. In Thm. C.19 we were able to deduce that $\theta_{a,a}=0\pmod{2\pi}$ by working with the topological spin rather than braiding. We now repeat this argument. Suppose that $a\in\mathrm{Ker}(g)$, i.e. that $\mathcal{T}a=a$ (modulo $c$). Then we will have that
@@
We note that the semion-fermion fermionic Abelian TO is well known to be anomalous Fidkowski et al. (2013); Wang and Senthil (2014). Even more clearly, if we impose $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ symmetry then we know that $Q_{a}=Q_{\mathcal{T}a}=Q_{a}+Q_{c}=Q_{a}+1\pmod{2}$, which is a contradiction. Thus we conclude that every fermionic Abelian TO with symmetry group $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ must have size $|\mathcal{A}|=2l^{2}$ for some $l$. We will nonetheless see that the other possibility will be useful when we generalize some results to the non-Abelian case.
@@
Let $\mathcal{A}$ be a time-reversal invariant Abelian bosonic TO. Suppose that it has an anyon $a\in\mathcal{A}$ which has order $n$, i.e. $n$ is the smallest integer such that $a^{n}=1$. Then we will have that
@@
We first establish something simple. Suppose that $1=(\mathcal{T}a)^{m}=\mathcal{T}a^{m}$, then we can use the fact that $\mathcal{T}^{2}$ is the identity and $\mathcal{T}1=1$ to see that $a^{m}=1$. Thus $m$ must be a multiple of $n$. We further note that $(\mathcal{T}a)^{n}=\mathcal{T}a^{n}=\mathcal{T}1=1$. We may then conclude that the order of $\mathcal{T}a$ is exactly $n$, i.e. equal to the order of $a$.
@@
Suppose that $0=\theta_{a,\mathcal{T}a}$. This will mean that $\theta_{a^{n},\mathcal{T}a^{m}}=nm\theta_{a,\mathcal{T}a}=0$, so every power of $a$ will braid trivially with every power of $\mathcal{T}a$. Then if $\mathcal{S}=\{1,a,\ldots,a^{n-1}\}\subseteq\mathcal{A}$ is the subgroup of powers of $a$, we see that =$\mathcal{R}=\{1,\mathcal{T}a,\ldots,\mathcal{T}a^{n-1}\}\subseteq\mathcal{S}^{\perp}$. Moreover, since $\mathcal{R}$ is a subgroup of $\mathcal{S}^{\perp}$ we know that $|\mathcal{S}^{\perp}|$ must be a multiple of $|\mathcal{R}|$ by Lagrange’s theorem, i.e. $|\mathcal{S}^{\perp}|=mn$ for some nonzero integer $m$. By Thm. 21 we then have that:
@@
Now suppose that $\theta_{a,\mathcal{T}a}=\pi$. Then we must have that $\theta_{a,\mathcal{T}a^{2}}=2\theta_{a,\mathcal{T}a}=0$, so $a$ braids trivially with $\mathcal{T}a^{2}$. Next suppose, by contradiction, that $n$ is odd, then $n+1$ is even and $(n+1)/2$ is an integer. Then we can see that $0=(n+1)/2\cdot\theta_{a,\mathcal{T}a^{2}}=\theta_{a,\mathcal{T}a^{n+1}}=\theta_{a,\mathcal{T}a}$ since $\mathcal{T}a^{n}=1$. This contradicts $\theta_{a,\mathcal{T}a}=\pi$, however. Thus if $\theta_{a,\mathcal{T}a}=\pi$ then $n$ is even. Then the set =$\mathcal{R}=\{1,\mathcal{T}a^{2},\mathcal{T}a^{4},\ldots,\mathcal{T}a^{n-2}\}$ is a subgroup of $\mathcal{A}$ with $|\mathcal{R}|=n/2$. Further if we take $\mathcal{S}=\{1,a,\ldots,a^{n-1}\}\subseteq\mathcal{A}$ then we can see that $\mathcal{R}\subseteq\mathcal{S}^{\perp}$. Since $\mathcal{R}$ is a subgroup of the group $\mathcal{S}^{\perp}$ this tells us that $|\mathcal{S}^{\perp}|$ must be a multiple of $n/2$ by Lagrange’s theorem, i.e. $|\mathcal{S}^{\perp}|=mn/2$ for some nonzero integer $m$. By Thm. 21 we have that $|\mathcal{A}|=mn^{2}/2$. Then by Thm. C.19 there must be some integer $l$ such that $l^{2}=mn^{2}/2$. Since $n$ is even we can write $n=2n^{\prime}$ for some integer $n^{\prime}$. Then we see that $l^{2}=2mn^{\prime 2}$ and thus $\sqrt{2m}=l/n^{\prime}$. Since the only integers whose square root is a rational number are the squares we know that $2m=k^{\prime 2}$ for some integer $k^{\prime}$. Then we see that $k^{\prime 2}$ is even and thus $k^{\prime}$ is even. Thus $k^{\prime}=2k$ for some integer $k$ and $m=2k^{2}$. We thus have $|\mathcal{A}|=mn^{2}/2=(nk)^{2}$. ∎
@@
Consider $\mathbb{Z}_{n}$ gauge theory. We know that it is time reversal invariant under the standard action of time reversal. Further we know that there exists $e\in\mathcal{A}$ which has order $n$. We also know that $|\mathcal{A}|=n^{2}$. Thus there is at least one time reversal invariant theory which saturates the lower bound given by Lem. C.23.
@@
Let $\mathcal{A}$ be a time-reversal invariant fermionic Abelian TO. Suppose it has an anyon $a\in\mathcal{A}$ with order $n$ such that $a^{k}\neq c$ for any $k$. Then we will have that
@@
First we note that since $a^{k}\neq c$ for any $k$ then $\mathcal{S}=\{1,a,\ldots,a^{n-1}\}$ forms a subgroup of $\mathcal{A}_{b}=\mathcal{A}/\{1,c\}$ of size $n$. We will work in this quotient for the rest of the proof. Recall that topological spin is then only defined modulo $\pi$, but braiding is unchanged. Now we must have that:
@@ C.3 $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ conserving fermionic Abelian TOs
We review some powerful results about $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ conserving fermionic Abelian TOs that were given in Lapa and Levin (2019) and conclude by extending them slightly. First we state a very useful theorem about fermionic Abelian TOs.
@@
The proof of this is given in Corollary A.19 of Drinfeld et al. (2010) and discussed in Lapa and Levin (2019). ∎
@@
This can be demonstrated with a simple example. Consider $\mathrm{U}(1)_{q}$, where $q$ is odd. We can express this as $\mathcal{A}=\{1,a,\ldots,a^{2q-1}\}$, where $\theta_{a}=\pi/q$. We see that $\theta_{a^{q},a^{n}}=2nq\theta_{a}=2n\pi=0\pmod{2\pi}$ and $\theta_{a^{q}}=q^{2}\theta_{a}=q\pi=\pi\pmod{2\pi}$ since $q$ is odd. So $a^{q}=c$ is a transparent fermion. Thus we can write $\mathcal{A}=\{1,a^{2},a^{4},\ldots,a^{2q-2}\}\times\{1,c\}$. The even powers of $a$ are closed under fusion and braiding is non-degenerate on this set.
@@
We repeat the arguments of Lapa and Levin (2019). Since $\mathrm{U}(1)$ charge is defined modulo two in this fermionic theory, then we must have that $\pi Q_{a}+\pi Q_{b}=\pi Q_{d}\pmod{2\pi}$ if $ab=d$. If we restrict to the bosonic TO $\mathcal{A}_{b}$, then we can use arguments about the non-degeneracy of the $S$-matrix Barkeshli et al. (2019) (see Eq. 45) to see that there must be a unique $\gamma\in\mathcal{A}_{b}$ such that $\theta_{\gamma,b}=\pi Q_{b}$ for every $b\in\mathcal{A}_{b}$.
@@
We then have that $\theta_{\gamma^{2},b}=2\pi Q_{b}$ for all $b\in\mathcal{A}$. Thus $\gamma^{2}$ has identical braiding to the vison, and since braiding is non-degenerate up to fusion with $c$ we must have that $\gamma^{2}=v$ or $vc$. ∎
@@
Consider $\mathbb{Z}_{2}$ gauge theory stacked with the fundamental fermion. Every anyon in this theory will square to the vacuum. Since the vacuum has even charge, this means every particle will have integer charge. In particular, the vison must braid trivially with every particle, so $v=1$. Given the natural charge assignment of the theory, we will have $\gamma=m$. We see that $\gamma$ has integer charge and $\gamma^{2}=v$ as it must.

### v2

@@ C.2 Time reversal invariant Abelian TOs
and thus $0=2\theta_{a}=\theta_{a,a}$ modulo $2\pi$. We thus have that $\theta_{a,a}=0$ for every $a\in\mathcal{C}=\mathrm{Ker}(g)/\mathrm{Im}(f)$. Now we note that $02\theta_{a,b}=\theta_{a^{2},b}=$ 0 since $a\in\mathcal{C}$ has order two. [^21] Note that here we are working in $\mathcal{C}$ so $a^{2}=f(a)\sim 1$. Thus $\theta_{a,b}=0,\pi$ for every $a,b\in\mathcal{C}$.
Now take some $a\in\mathcal{C}$, then $\theta_{a,a}=0$ as demonstrated above. Since braiding is nondegenerate on $\mathcal{C}$ there must be some $b\in\mathcal{C}$ such that $\theta_{a,b}=\pi$. Then take $\mathcal{S}=\{1,a,b,ab\}\subseteq\mathcal{C}$. Then by Thm. 21 we see that $|\mathcal{C}|=|\mathcal{S}||\mathcal{S}^{\perp}|=4|\mathcal{S}^{\perp}|$. We further know by Corr. C.9 that $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$, so the only elements which braid trivially with everything in $\mathcal{S}^{\perp}$ lie in $\mathcal{S}$. But every nontrivial element of $\mathcal{S}$ braids nontrivially with at least one other element of $\mathcal{S}$ and thus $\mathcal{S}\cap\mathcal{S}^{\perp}=\{1\}$. We conclude that braiding is nondegenerate when restricted to $\mathcal{S}^{\perp}$. Thus $\mathcal{S}^{\perp}\subseteq\mathcal{C}$ is itself a bosonic TO. We can then find some $\mathcal{S}^{\prime}=\{1,a^{\prime},b^{\prime},a^{\prime}b^{\prime}\}\subseteq\mathcal{S}^{\perp}$ where $\theta_{a^{\prime},b^{\prime}}=\pi$ and again apply Thm. 21 to see that $|\mathcal{S}^{\perp}|=4|\mathcal{S}^{\prime\perp}|$, where $\mathcal{S}^{\prime\perp}\subseteq\mathcal{S}^{\perp}$ is another bosonic TO. Continuing to apply this logic until we have exhausted the finite set $\mathcal{C}$ reveals that $|\mathcal{C}|=4^{m}=2^{2m}$ for some $m$. Thus $|\mathcal{C}|$ is a square and we see $|\mathcal{A}_{b}|=|\mathrm{Im}(f)|^{2}|\mathcal{C}|$ must also be a square. ∎
@@
If $\mathcal{A}$ is a time-reversal invariant fermionic Abelian TO then we write $\mathcal{A}_{b}=\mathcal{A}/\{1,c\}$. Braiding will become nondegenerate on this TO. This is thus a bosonic TO with a $\pi$-ambiguity in the topological spin, since we have identified the fundamental fermion $c$ with the vacuum. However, there is no ambiguity of braiding.
We can now repeat the initial arguments of Thm. C.19 to see that $|\mathcal{A}_{b}|=|\mathrm{Im}(f)|^{2}|\mathcal{C}|$, where $\mathcal{C}=\mathrm{Ker}(g)/\mathrm{Im}(f)$. We continue to have that every element of $\mathcal{C}$ is order two, so $\mathcal{C}=\mathbb{Z}_{2}^{n}$ for some $n$. These results did not rely on topological spin, only on braiding, so they continue to hold in $\mathcal{A}_{b}$. Then we know that $|\mathcal{A}|=2|\mathcal{A}_{b}|=|\mathrm{Im}(f)|^{2}2^{n+1}$. If $n$ is odd then $|\mathcal{A}|=l^{2}$ for some $l$, while if $n$ is even then $|\mathcal{A}|=2l^{2}$ for some $l$.
We need to work a little harder to prove the second result since Eq. (33) did rely on topological spin. We now revisit the argument surrounding this equation. Suppose that $a\in\mathrm{Ker}(g)$ i.e., that $\mathcal{T}a=a$ (modulo $c$). Then we have that
@@
We note that the semion-fermion fermionic Abelian TO is well known to be anomalous Fidkowski et al. (2013); Wang and Senthil (2014). Even more clearly, if we impose $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ symmetry then we know that $Q_{a}=Q_{\mathcal{T}a}=Q_{a}+Q_{c}=Q_{a}+1\pmod{2}$, which is a contradiction. Thus we conclude that every fermionic Abelian TO with symmetry group $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ must have size $|\mathcal{A}|=2l^{2}$ for some $l$. We nonetheless see that the other possibility will be useful when we generalize some results to the non-Abelian case.
@@
Let $\mathcal{A}$ be a time-reversal invariant Abelian bosonic TO. Suppose that it has an anyon $a\in\mathcal{A}$ which has order $n$ i.e., $n$ is the smallest integer such that $a^{n}=1$. Then we have that
@@
We first establish something simple. Suppose that $1=(\mathcal{T}a)^{m}=\mathcal{T}a^{m}$, then we can use the fact that $\mathcal{T}^{2}$ is the identity and $\mathcal{T}1=1$ to see that $a^{m}=1$. Thus $m$ must be a multiple of $n$. We further note that $(\mathcal{T}a)^{n}=\mathcal{T}a^{n}=\mathcal{T}1=1$. We may then conclude that the order of $\mathcal{T}a$ is exactly $n$ i.e., equal to the order of $a$.
@@
Suppose that $0=\theta_{a,\mathcal{T}a}$. This will mean that $\theta_{a^{n},\mathcal{T}a^{m}}=nm\theta_{a,\mathcal{T}a}=0$, so every power of $a$ will braid trivially with every power of $\mathcal{T}a$. Then if $\mathcal{S}=\{1,a,\ldots,a^{n-1}\}\subseteq\mathcal{A}$ is the subgroup of powers of $a$, we see that =$\mathcal{R}=\{1,\mathcal{T}a,\ldots,\mathcal{T}a^{n-1}\}\subseteq\mathcal{S}^{\perp}$. Moreover, since $\mathcal{R}$ is a subgroup of $\mathcal{S}^{\perp}$ we know that $|\mathcal{S}^{\perp}|$ must be a multiple of $|\mathcal{R}|$ by Lagrange’s theorem i.e., $|\mathcal{S}^{\perp}|=mn$ for some nonzero integer $m$. By Thm. 21 we then have that:
@@
Now suppose that $\theta_{a,\mathcal{T}a}=\pi$. Then we must have that $\theta_{a,\mathcal{T}a^{2}}=2\theta_{a,\mathcal{T}a}=0$, so $a$ braids trivially with $\mathcal{T}a^{2}$. Next suppose, by contradiction, that $n$ is odd, then $n+1$ is even and $(n+1)/2$ is an integer. Then we can see that $0=[(n+1)/2]\theta_{a,\mathcal{T}a^{2}}=\theta_{a,\mathcal{T}a^{n+1}}=\theta_{a,\mathcal{T}a}$ since $\mathcal{T}a^{n}=1$. This contradicts $\theta_{a,\mathcal{T}a}=\pi$, however. Thus if $\theta_{a,\mathcal{T}a}=\pi$ then $n$ is even. Then the set =$\mathcal{R}=\{1,\mathcal{T}a^{2},\mathcal{T}a^{4},\ldots,\mathcal{T}a^{n-2}\}$ is a subgroup of $\mathcal{A}$ with $|\mathcal{R}|=n/2$. Further if we take $\mathcal{S}=\{1,a,\ldots,a^{n-1}\}\subseteq\mathcal{A}$ then we can see that $\mathcal{R}\subseteq\mathcal{S}^{\perp}$. Since $\mathcal{R}$ is a subgroup of the group $\mathcal{S}^{\perp}$ this tells us that $|\mathcal{S}^{\perp}|$ must be a multiple of $n/2$ by Lagrange’s theorem i.e., $|\mathcal{S}^{\perp}|=mn/2$ for some nonzero integer $m$. By Thm. 21 we have that $|\mathcal{A}|=mn^{2}/2$. Then by Thm. C.19 there must be some integer $l$ such that $l^{2}=mn^{2}/2$. Since $n$ is even we can write $n=2n^{\prime}$ for some integer $n^{\prime}$. Then we see that $l^{2}=2mn^{\prime 2}$ and thus $\sqrt{2m}=l/n^{\prime}$. Since the only integers whose square root is a rational number are the squares we know that $2m=k^{\prime 2}$ for some integer $k^{\prime}$. Then we see that $k^{\prime 2}$ is even and thus $k^{\prime}$ is even. Thus $k^{\prime}=2k$ for some integer $k$ and $m=2k^{2}$. We thus have $|\mathcal{A}|=mn^{2}/2=(nk)^{2}$. ∎
@@
Consider $\mathbb{Z}_{n}$ gauge theory. We know that it is time-reversal invariant under the standard action of time reversal. Further we know that there exists $e\in\mathcal{A}$ which has order $n$. We also know that $|\mathcal{A}|=n^{2}$. Thus there is at least one time reversal invariant theory which saturates the lower bound given by Lem. C.23.
@@
Let $\mathcal{A}$ be a time-reversal invariant fermionic Abelian TO. Suppose it has an anyon $a\in\mathcal{A}$ with order $n$ such that $a^{k}\neq c$ for any $k$. Then we have that
@@
First we note that since $a^{k}\neq c$ for any $k$ then $\mathcal{S}=\{1,a,\ldots,a^{n-1}\}$ forms a subgroup of $\mathcal{A}_{b}=\mathcal{A}/\{1,c\}$ of size $n$. We work in this quotient for the rest of the proof. Recall that topological spin is then only defined modulo $\pi$, but braiding is unchanged. Now we must have that:
@@ C.3 $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ conserving fermionic Abelian TOs
We review some powerful results about $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ conserving fermionic Abelian TOs that were given in Ref. Lapa and Levin (2019) and conclude by extending them slightly. First we state a very useful theorem about fermionic Abelian TOs.
@@
The proof of this is given in Corollary A.19 of Ref. Drinfeld et al. (2010) and discussed in Ref. Lapa and Levin (2019). ∎
@@
This can be demonstrated with a simple example. Consider $\mathrm{U}(1)_{q}$, where $q$ is odd. We can express this as $\mathcal{A}=\{1,a,\ldots,a^{2q-1}\}$, where $\theta_{a}=\pi/q$. We see that $\theta_{a^{q},a^{n}}=2nq\theta_{a}=2n\pi=0\pmod{2\pi}$ and $\theta_{a^{q}}=q^{2}\theta_{a}=q\pi=\pi\pmod{2\pi}$ since $q$ is odd. So $a^{q}=c$ is a transparent fermion. Thus we can write $\mathcal{A}=\{1,a^{2},a^{4},\ldots,a^{2q-2}\}\times\{1,c\}$. The even powers of $a$ are closed under fusion and braiding is nondegenerate on this set.
@@
We repeat the arguments of Ref. Lapa and Levin (2019). Since $\mathrm{U}(1)$ charge is defined modulo two in this fermionic theory, then we must have that $\pi Q_{a}+\pi Q_{b}=\pi Q_{d}\pmod{2\pi}$ if $ab=d$. If we restrict to the bosonic TO $\mathcal{A}_{b}$, then we can use arguments about the nondegeneracy of the $S$-matrix Barkeshli et al. (2019) [see Eq. (45)] to see that there must be a unique $\gamma\in\mathcal{A}_{b}$ such that $\theta_{\gamma,b}=\pi Q_{b}$ for every $b\in\mathcal{A}_{b}$.
@@
We then have that $\theta_{\gamma^{2},b}=2\pi Q_{b}$ for all $b\in\mathcal{A}$. Thus $\gamma^{2}$ has identical braiding to the vison, and since braiding is nondegenerate up to fusion with $c$ we must have that $\gamma^{2}=v$ or $vc$. ∎
@@
Consider $\mathbb{Z}_{2}$ gauge theory stacked with the fundamental fermion. Every anyon in this theory will square to the vacuum. Since the vacuum has even charge, this means every particle will have integer charge. In particular, the vison must braid trivially with every particle, so $v=1$. Given the natural charge assignment of the theory, we have $\gamma=m$. We see that $\gamma$ has integer charge and $\gamma^{2}=v$ as it must.